# Asymptotic spreading of predator-prey populations in a shifting environment

### v1

@@ 6 Convergence to homogeneous state
In this section, we apply the previous spreading result to characterize the long-term behavior of solutions of (4) in the moving frame where the predator persists.
Having established the existence of the spreading speed $c_{*}$, it follows that $(u,v)\to(0,1)$ locally uniformly in any moving frames with speed above $c_{*}$, and that $u$ persists in any moving frames with speed below $c_{*}$. In the following we discuss the asymptotic behavior of the soltuions in the latter case. Define, for $i=1,2$,
@@
$(\tilde{u},\tilde{v})=\begin{cases}(u_{2},v_{2})&\text{ if }c_{1}<c^{*}\quad\text{ and }x_{k}\in\big{(}(c_{1}+\eta)t,(c^{*}-\eta)t\big{)},\\ (u_{1},v_{1})&\text{ otherwise.}\end{cases}$ (69)
@@
Indeed, given $K$, choose $R>0$ such that $K\subset[-R,R]\times[-R,R]$, then for $k\gg 1$, we have
$K\subset[-R,R]^{2}$$\subseteq\{(t,x):|t|\leq R,\quad\eta(R-t_{k})\leq x\leq-(c^{*}-\eta/2)R+\eta t_{k}/2\}$
$\subseteq\{(t,x):\leavevmode\nobreak\ -\eta(t+t_{k})\leq x\leq(c^{*}-\eta/2)t+\eta t_{k}/2\}$
$\subseteq\{(t,x):\leavevmode\nobreak\ -\eta(t+t_{k})-x_{k}\leq x\leq(c^{*}-\eta/2)t-x_{k}+(c^{*}-\eta/2)t_{k}\}=\Omega_{k}.$
@@
Using the above $L^{\infty}$ bounds and parabolic $L^{p}$ estimates we may deduce, re-labelling a sub-sequence if necessary, that $(u_{k},v_{k})$ weakly in $W^{2,1,p}_{loc}(\mathbb{R}^{2})$ (and strongly in $C^{1+\alpha,(1+\alpha)/2}_{loc}(\mathbb{R}^{2})$ thanks to Sobolev embedding) to an entire solution $(u_{\infty},v_{\infty})$ of the system
@@
where $\tilde{a}=a_{2}$ if $c_{1}<c^{*}$ and $x_{k}\in\big{(}(c_{1}+\eta)t,(c^{*}-\eta)t\big{)}$, and $\tilde{a}=a_{1}$ otherwise. Moreover, (70) also implies that
@@ Appendix A Comparison Principle
The assumptions of $\tilde{R}$ is stated precisely as follows.
@@
[…]
@@
We apply the results of [40]. The specific form of the Hamilton-Jacobi equation (72) and assumptions (B1)-(B2), imply that [40, (H1)-(H6)] hold. Hence, [40, Proposition 2.11] applies. ∎
@@
Thanks to (B1)-(B2), [40, Proposition 1.7(b)] applies. Next, by applying [40, Lemma 2.9], we deduce that $s\mapsto\hat{\rho}(s)$ is nondecreasing. ∎

### v2

@@ 6 Convergence to homogeneous state
In this section, we apply the previous spreading result to characterize the long-time behavior of solutions of (4) in the moving frame where the predator persists.
Having established the existence of the spreading speed $c^{*}$, and recalling Lemma 2, it follows that $(u,v)\to(0,1)$ locally uniformly in any moving frames with speed above $c^{*}$, and that $u$ persists in any moving frames with speed below $c^{*}$. In the following we discuss the asymptotic behavior of the solutions in the latter case. Define, for $i=1,2$,
@@
$(\tilde{u},\tilde{v})=\begin{cases}(u_{2},v_{2})&\text{ if }c_{1}<c^{*}\quad\text{ and }x_{k}\in\big{(}(c_{1}+\eta)t_{k},(c^{*}-\eta)t_{k}\big{)},\\ (u_{1},v_{1})&\text{ otherwise.}\end{cases}$ (60)
@@
Indeed, given $K$, choose $R>0$ such that $K\subset[-R,R]\times[-R,R]$. Then for $k\gg 1$, we have
$K$ $\subset[-R,R]^{2}$
$\subseteq\{(t,x):|t|\leq R,\quad\eta(R-t_{k})\leq x\leq-(c^{*}-\eta/2)R+\eta t_{k}/2\}$
$\subseteq\{(t,x):|t|\leq R,\quad-\eta(t+t_{k})\leq x\leq(c^{*}-\eta/2)t+\eta t_{k}/2\}$
$\subseteq\{(t,x):t+t_{k}\geq T,\quad-\eta(t+t_{k})-x_{k}\leq x\leq(c^{*}-\eta/2)t-x_{k}+(c^{*}-\eta/2)t_{k}\}$
$=\Omega_{k}.$
@@
Using the above $L^{\infty}$ bounds and parabolic $L^{p}$ estimates we may deduce, re-labelling a sub-sequence if necessary, that $(u_{k},v_{k})$ converges weakly in $W^{2,1,p}_{loc}(\mathbb{R}^{2})$ (and strongly in
$C^{1+\alpha,(1+\alpha)/2}_{loc}(\mathbb{R}^{2})$ thanks to Sobolev embedding) to an entire solution $(u_{\infty},v_{\infty})$ of the system
@@
where $\tilde{a}=a_{2}$ if $c_{1}<c^{*}$ and $x_{k}\in\big{(}(c_{1}+\eta)t_{k},(c^{*}-\eta)t_{k}\big{)}$, and $\tilde{a}=a_{1}$ otherwise. Moreover, (61) also implies that
@@ Appendix A Comparison Principle
The assumptions on $\tilde{R}$ are stated precisely as follows.
@@
$\underline{\rho}(0)\leq\overline{\rho}(0)\quad\text{and}\quad\lim_{s\to\infty}\frac{\underline{\rho}(s)}{s}\leq\lim_{s\to\infty}\frac{\overline{\rho}(s)}{s}.$ (64)
@@
We apply the results of [41]. The specific form of the Hamilton-Jacobi equation (63) and assumptions (B1)-(B2), imply that [41, (H1)-(H6)] hold. Hence, [41, Proposition 2.11] applies. ∎
@@
Thanks to (B1)-(B2), [41, Proposition 1.7(b)] applies. Next, by applying [41, Lemma 2.9], we deduce that $s\mapsto\hat{\rho}(s)$ is nondecreasing. ∎